OpenStudy (anonymous):

Consider the numbers 4/13 and 3/10 a. Compare the numbers. b. Find a rational number between the two given numbers Please help

OpenStudy (anonymous):

It says I have to show work but I have no idea how to do this problem

OpenStudy (aaliyahfairgood):

find a common denominator which means you find the least common multiple in the factors 10 and 13 4/13 * 10/10 = 40/130 3/10 * 13/13= 39/130

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Then what?

OpenStudy (anonymous):

Helloooo?

OpenStudy (aaliyahfairgood):

omg srry I was afk

OpenStudy (anonymous):

It's cool

OpenStudy (aaliyahfairgood):

ok so you look at the numerators and since 3/10 is less than 4/13 when you look at it as 39/130 is less than 40/130, then the numerator ratio is

OpenStudy (aaliyahfairgood):

39 to 40

OpenStudy (aaliyahfairgood):

looking at the numerators

OpenStudy (anonymous):

Well then what is the rational number between two of the numbers?